# Scaling Diffusion Transformers to 16 Billion Parameters

### v1

@@ 3.1 Experimental Settings
We use the AdamW optimizer [50] without weight decay across all datasets, maintaining a constant learning rate of 1e-4. In line with [68], we utilize an exponential moving average of DiT-MoE weights over training with a decay of 0.9999. All results were reported using the EMA model. Our models are trained on Nvidia A100 GPU. When trained on ImageNet dataset at different resolutions, we adopt classifier-free guidance [41] following [78] and use an off-the-shelf pre-trained variational autoencoder (VAE) model [52] from Stable Diffusion [78] available in huggingface [^1] https://huggingface.co/stabilityai/sd-vae-ft-ema. The VAE encoder has a downsampling factor of 8. We retrain diffusion hyper-parameters from [68], using a $t_{max}=1000$ linear variance schedule ranging from $1\times 10^{-4}$ to $2\times 10^{-2}$ and parameterization of the covariance. We set the share experts number $n_{s}$ to 2 and the expert-level balance factor $\alpha$ to 0.05 by default.
@@
Figure 4: Training loss curves for small version variants. We can see that the incorporation of simple designs helps DiT-MoE training stably and consistently converge. On the other hand, increasing the expert number helps loss decrease while introducing more loss spikes.
@@ 3.2 Model Design Analysis
We directly increase the expert number from 8 to 16, while keeping the number of activated parameters fixed at 2. As reported in Figure 3 (b), the adjustment leads to consistently improved generative performance, albeit with a significant increase in GPU memory consumption. On the other hand, the loss curve in Figure 4 demonstrates that incorporation of MoE can achieve competitive performance and helps to faster loss convergence. However, directly increasing the expert number for performance enhancement may introduce more loss spikes. We leave how to reduce loss spike, such as spike regularization loss [98], in future work.
@@ 4 Related Works
While previous studies predominantly utilize a single model to tackle denoising tasks across various timesteps [68, 32, 69, 31, 28, 46], several investigations have explored the deployment of multiple expert models, each specializing in a distinct range of timesteps [13]. PPAP [36] achieves this by training multiple classifiers on segmented timesteps, each employed for classifier guidance. Similarly, e-DiffI [1] and ERNIE-ViLG [33] utilize a consistent set of denoisers across these experts, whereas MEME [54] advocates for distinct architectures tailored to each timestep segment. These methodologies enhance generative quality while maintaining comparable inference costs, albeit at the expense of increased memory requirements. They operate on the premise that the characteristics of denoising tasks vary significantly across timesteps. We extend it by analyzing the expert routing mechanism and demonstrating that both temporal and spatial elements without class-conditional information influence different MoE layers. The most similar to work is [64], which also explores experts routing, however, they focus on a form of multi-task learning for time step and not actually sparse, i.e., base version vs. dense version comes to 144M vs. 131M. In contrast, we delve into the time-space routing mechanism and modeling of $>$10B model size.

### v2

@@ 3.1 Experimental Settings
We use the AdamW optimizer [50] without weight decay across all datasets, maintaining a constant learning rate of 1e-4. In line with [69], we utilize an exponential moving average of DiT-MoE weights over training with a decay of 0.9999. All results were reported using the EMA model. Our models are trained on Nvidia A100 GPU. When trained on ImageNet dataset at different resolutions, we adopt classifier-free guidance [41] following [79] and use an off-the-shelf pre-trained variational autoencoder (VAE) model [52] from Stable Diffusion [79] available in huggingface [^1] https://huggingface.co/stabilityai/sd-vae-ft-ema. The VAE encoder has a downsampling factor of 8. We retrain diffusion hyper-parameters from [69], using a $t_{max}=1000$ linear variance schedule ranging from $1\times 10^{-4}$ to $2\times 10^{-2}$ and parameterization of the covariance. We set the share experts number $n_{s}$ to 2 and the expert-level balance factor $\alpha$ to 0.005 by default.
@@
Figure 4: Training loss curves for in small version for different elements. (a) The effect of expert balance loss. We can see that 0.005 can achieve best performance. (b) The effect of experts number with same activate experts. We can see that increase expert number results consistently performance enhancement. (c) The influence of MoE layer replacement. If only half of MLP are replace with MoE layer, we can see that MoE in deep layer results in better performance, which is consisent with experted in visualization. Note that Figures are experimented in rectifide flow training.
@@ 3.2 Model Design Analysis
Effect of expert-level balance loss.
As a import part of balance expert loading, we range from 0.5, 0.01 to 0.05, 0.005. We also dropout the expert balance loss, i.e., set to zero. As piloted in Figure 4 (a), as we expert, increase the expert-level balance loss leads to a performance first improve then decrease. The optimal factor is 0.005. Meantime, we leave how to dynamically adjust the factor according to current training situation in future work.
@@
We directly increase the expert number from 8 to 16, while keeping the number of activated parameters fixed at 2. As reported in Figure 3 (b), the adjustment leads to consistently improved generative performance, albeit with a significant increase in GPU memory consumption. On the other hand, the loss curve in Figure 4(b) also demonstrates that incorporation of MoE in text to image tasks can achieve competitive performance and helps to faster loss convergence.
@@ 3.3 Expert Specialization Analysis
Base on the above observation, we assume that the shallow layer in DiT is sparse, and more capacity is required in deep layer. Therefore, we try to replace the dense DiT with half of MoE layer. Specifically, we replace with shallow half ($0\sim\frac{N}{2}$) and deep half ($\frac{N}{2}\sim N$) layers with MoE layers. The results in Figure 4(c) show that replacing MoE in deep layers achieve lower loss and better generative performance.
@@ 4 Related Works
While previous studies predominantly utilize a single model to tackle denoising tasks across various timesteps [69, 32, 70, 31, 28, 46], several investigations have explored the deployment of multiple expert models, each specializing in a distinct range of timesteps [13]. PPAP [36] achieves this by training multiple classifiers on segmented timesteps, each employed for classifier guidance. e-DiffI [1] and ERNIE-ViLG [33] utilize a consistent set of denoisers across these experts, whereas MEME [54] advocates for distinct architectures tailored to each timestep segment. Additionally, [64] deals with routing problem of denoising tasks across timesteps and designs a routing strategies for learning these tasks in a single model. These methodologies enhance generative quality while maintaining comparable inference costs, albeit at the expense of increased memory requirements. They operate on the premise that the characteristics of denoising tasks vary across timesteps. We extend it by analyzing the expert routing mechanism and demonstrating that both temporal and spatial elements without class-conditional information influence different MoE layers. The most similar to work is [65], which also explores experts routing, however, they focus on a form of multi-task learning for time step and not actually sparse, i.e., base version vs. dense version comes to 144M vs. 131M. In contrast, we delve into the time-space routing mechanism and modeling of $>$10B model size.